g(x)=x-1/x f(g(x))=x^3-1/x^3 then prove that f(x)=3x^2+3

g(x)=x-1/x   f(g(x))=x^3-1/x^3


then prove that f(x)=3x^2+3

Grade:11

1 Answers

pankaj dave
11 years ago

f(g(x))=(x-1/x)3+3(x.1/x)(x-1/x)

 f(y)  =y3+3y        y=x-1/x